Queens Soliderate is considering two mutually exclusive projects. The firm, which has a 12% cost of capital, has estimated its cash flows as shown in the following table.

Project A Project B

Initial investment (CF0) $130,000 $85,000

Year (t) Cash inflows (CFt)

1 $25,000 $40,000

2 35,000 35,000

3 45,000 30,000

4 50,000 10,000

5 55,000 5,000

a. Calculate the NPV of each project, and assess its acceptability. (1.5)

b. Calculate the IRR for each project, and assess its acceptability. (1.5)Required to answer. Single line text.

(3 Points)


1
Expert's answer
2021-06-22T17:42:21-0400

(a) 

a. Project A

  "NPV = F \div [ (1 + r)^n ]"

where, PV = Present Value, F = Future payment (cash flow), r = Discount rate, n = the number of periods in the future.







NPV= cash inflows-cash outflows

   = 145245-130000

   =15245

Project B

"NPV = F \div [ (1 + r)^n ]"

where, PV = Present Value, F = Future payment (cash flow), r = Discount rate, n = the number of periods in the future.







NPV=Cash inflows-cash outflows

    =94170-85000

    =9170

Project A will be selected because it is having a higher NPV than Project1 B. 


(b) Internal rate of return (IRR) = The rate of return where Present value of cash inflows equals the Present value of cash outflows

"Project A, IRR\\: \$0 = \$25,000\div(1+IRR)^1 + \$35,000\div(1+IRR)^2 + \$45,000\div(1+IRR)^3 + \$50,000\div(1+IRR)^4 + \$55,000\div(1+IRR)^5 - \$130,000 \\IRR = 16\%"

Project A IRR = 16%, the project is acceptable since the IRR is greater than the 12% cost of capital

"Project B, IRR\\: \$0 = \$40,000\div(1+IRR)^1 + \$35,000\div(1+IRR)^2 + \$30,000\div(1+IRR)^3 + \$10,000\div(1+IRR)^4 + \$5,000\div(1+IRR)^5 - \$85,000\\ IRR = 18\%"

Project B IRR = 18%, the project is acceptable since the IRR is greater than the 12% cost of capital